the sum of the measures of the interior angles of a polygon is the same as the sum of the measures of its exterior angles. how many sides does the polygon have?
Answer by Don2xmalabag(6)

(n-2)180=360: where n=number of sides
{(n-2)180}/180=360/180
n-2=2
n-2+2=2+2
n=4
*therefore the polygon is a Quadrilateral